OpenStudy (anonymous):

PLEASE HELP!!!! What is the solution to the following system of equations: 3x + 4y + z = 10 2x - 5z = 4 y = 1

jimthompson5910 (jim_thompson5910):

3x + 4y + z = 10 3x + 4(1) + z = 10 3x + 4 + z = 10 3x + z = 10-4 3x + z = 6 So we have two equations with two unknowns (x and z) 3x + z = 6 2x - 5z = 4 Solve the first equation for z 3x + z = 6 z = 6 - 3x and then plug in z = 6 - 3x into the second equation 2x - 5z = 4 2x - 5(6-3x) = 4 2x - 30+15x = 4 Now solve for x 2x - 30+15x = 4 17x - 30 = 4 17x = 4+30 17x = 34 x = 34/17 x = 2 Now that we know that x = 2, we can use this to find z z = 6 - 3x z = 6 - 3(2) z = 6 - 6 z = 0 So the solutions are x = 2, y = 1, and z = 0 They can be represented as the ordered triple (2,1,0)

OpenStudy (anonymous):

Wow! Look at you go, again! Thank you so much - I really do appreciate your thoroughness. I know how long that takes to type and keep your brain and fingers straight, at the same time! :)

jimthompson5910 (jim_thompson5910):

lol it's not so bad really

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

exactly my thoughts, i love doing this (weird I know huh lol)

OpenStudy (anonymous):

No, not at all! :) I love to help teach, as well (maybe just not so much, in the area of math). I don't think it's weird, at all. It's just inside some people.